OpenStudy (anonymous):

6x + 2 + 3x = 10x + 2 – x Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution)

OpenStudy (anonymous):

Bring x terms to LHS, numbers to RHS: 6x + 3x - 10x +x = 2 - 2 = 0 So 10x - 10x = 0 = 0, so null equation (no solutions)